OpenStudy (anonymous):

4x2 - 16 difference of squares

jimthompson5910 (jim_thompson5910):

Rewrite 4x^2 into (2x)^2. Rewrite 16 as 4^2 So 4x^2-16 turns into (2x)^2-4^2. So you can see we have a difference of squares. Now use the idea that a^2-b^2 = (a-b)(a+b). In this case, a = 2x and b = 4

OpenStudy (anonymous):

no im stating its a difference of squares am i correct?

jimthompson5910 (jim_thompson5910):

oh yes, you are correct

jimthompson5910 (jim_thompson5910):

Each expression can be written as a square of something else. So they are both squares

OpenStudy (anonymous):

Miss, do you want factor expresion

OpenStudy (anonymous):

o ok yeah i wasnt asking for you to solve the problem thanks

jimthompson5910 (jim_thompson5910):

you're welcome